Find the area of the parallelogram with vertices
A(2, -3), B(7, -3), C(9, 2), D(4, 2)

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52786)

.
The base AB is a horizontal line and its length is 7 - 2 = 5 units.


The height is vertical line and its length is 2 - (-3) = 5.


Hence, the area of the parallelogram is 5*5 = 25 square units.


Answer by MathTherapy(10552)

Find the area of the parallelogram with vertices
A(2, -3), B(7, -3), C(9, 2), D(4, 2)
Lines AB and CD are horizontal, are parallel, and measure 5 units each. Either one can be the base of the parallelogram
The height, or perpendicular segment from D to base AB is 5 (2 - - 3) units
Hence, bh, or